# Krylov complexity of thermal state in early universe

### v1

@@ 1 Introduction
In recent years, the Krylov complexity has made rapid progress in condensed matter and high-energy physics. Ref. Muck:2022xfc has systematically investigated the Krylov complexity in various polynomials dubbed as the inner product. The Krylov complexity has been applied to the SKY model Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk , the generalised coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye . The chaos can be understood as a delocalization in the Krylov space Dymarsky:2019elm . Thereafter, Ref. Bhattacharjee:2022vlt has shown that the exponential growth of Krylov complexity can be observed in integrable systems with saddle-dominated scrambling, where Ref. Huh:2023jxt verified all the universal features of spread complexity in this framework. The very recent developments of Krylov complexity can be found in Refs. Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj . Especially, the detailed implication of Krylov complexity in various fields can be found in Ref. Nandy:2024htc .
@@ 4 The wave function of thermal state
Physically speaking, the Hamiltonian of single-field inflation can be considered as an inverted harmonic oscillator Li:2024kfm . Meanwhile, we will investigate the Krylov complexity for the thermal state.
@@
In the context of bosonic oscillators, where $\beta=\frac{1}{k_{B}T}$ and $E_{n}=n\omega$ ($\omega$ is the frequency), we will consider the thermal state (17) as a pure state following the approach in Jefferson:2017sdb ; Ali:2018fcz , using the method described in Haque:2021kdm . Given that the thermal state is a mixed state, we need to create the purification of $\hat{\rho}$, which we refer to as a pure state $|\Psi\rangle$, in an enlarged Hilbert space $\mathcal{H}_{pure}=\mathcal{H}\otimes\mathcal{H}_{anc}$, where $\mathcal{H}_{anc}$ represents the ”ancillary” degrees of freedom. When we trace $|\Psi\rangle$ over $\mathcal{H}_{anc}$, we obtain the mixed state $\hat{\rho}_{mix}$, indicating that $|\Psi\rangle$ serves as the purification of the thermal state (17).
@@ 5 Evolution of temperature and squeezed angle
With the application of Fourier decomposition,
@@
As a result, we could obtain the Hamiltonian in terms of creation and annihilation operators,
@@ 7.4 Some discussions for $\mu_{2}$
In our previous study Li:2024kfm , we already concluded that the inflation is a strong dissipative system leading to $\mu_{2}\gg 1$, and we also guessed that the Krylov complexity will saturate some constant values in RD and MD which the RD and MD may be the weak dissipative region. In this section, we will confirm this point. First, from the evolution trend of Krylov complexity in both methods, the method of closed system and the method of open system. From Figs. 4 and 6, it indicates that the Krylov complexity will saturate some constant values which is consistent with the discussions of Ref. Parker:2018yvk in the weak dissipative region. Another reason is from the explicit formula of $\mu_{2}=(\frac{a^{2}}{2k}V_{,\phi\phi}+k)$, if we take values of parameters that are given in Figs. 4 and 6, we could see that $\mu_{2}\ll 1$ in RD and MD. Consequently, our numeric has confirmed our previous guess which the RD and MD are the weak dissipative system.

### v2

@@ 1 Introduction
In recent years, the Krylov complexity has made rapid progress in condensed matter and high-energy physics. Ref. Muck:2022xfc has systematically investigated the Krylov complexity in various polynomials dubbed as the inner product. The Krylov complexity has been applied to the SKY model Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk , the generalised coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye . The chaos can be understood as a delocalization in the Krylov space Dymarsky:2019elm . Thereafter, Ref. Bhattacharjee:2022vlt has shown that the exponential growth of Krylov complexity can be observed in integrable systems with saddle-dominated scrambling, where Ref. Huh:2023jxt verified all the universal features of spread complexity in this framework. The very recent developments of Krylov complexity can be found in Refs. Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj ; Camargo:2023eev ; Camargo:2024deu ; Baggioli:2024wbz . Especially, the detailed implication of Krylov complexity in various fields can be found in Ref. Nandy:2024htc .
@@ 4 The wave function of thermal state
We will investigate the Krylov complexity for the thermal state.
@@
In the context of bosonic oscillators, where $\beta=\frac{1}{k_{B}T}$ and $E_{n}=n\omega$ ($\omega$ is the frequency), we will consider the thermal state (17) as a pure state following the approach in Jefferson:2017sdb ; Ali:2018fcz , using the method described in Haque:2021kdm . Since the thermal state is a mixed state, we need to create the purification of $\hat{\rho}$, which refers as a pure state $|\Psi\rangle$, in an enlarged Hilbert space $\mathcal{H}_{pure}=\mathcal{H}\otimes\mathcal{H}_{anc}$, where $\mathcal{H}_{anc}$ represents the ”ancillary” degrees of freedom. When we trace $|\Psi\rangle$ over $\mathcal{H}_{anc}$, we obtain the mixed state $\hat{\rho}_{mix}$, indicating that $|\Psi\rangle$ serves as the purification of the thermal state (17). Similar method can also be found in Das:2024zuu , which they can also map the pure state into the mixed state.
@@ 5 Evolution of temperature and squeezed angle
With the application of Fourier decomposition and Fourier transformation,
@@
As a result, we could obtain the Hamiltonian in terms of creation and annihilation operators in momentum space,
@@ 7.4 Some discussions for $\mu_{2}$
In our previous study Li:2024kfm , we already concluded that the inflation is a strong dissipative system leading to $\mu_{2}\gg 1$, and we also guessed that the Krylov complexity will saturate some constant values in RD and MD which the RD and MD may be the weak dissipative region. In this section, we will confirm this point. First, from the evolution trend of Krylov complexity in both methods, the method of closed system and the method of open system. From Figs. 4 and 6, it indicates that the Krylov complexity will saturate some constant values which is consistent with the discussions of Ref. Parker:2018yvk in the weak dissipative region. Another reason is from the explicit formula of $\mu_{2}=(\frac{a^{2}}{2k}V_{,\phi\phi}+k)$, if we take values of parameters that are given in Figs. 4 and 6, we could see that $\mu_{2}\ll 1$ in RD and MD. Physically speaking, the energy of inflationary potential will transfer into the other field that is dubbed as preheating, where this process will lead to the generation of particles, we could conclude that the process of preheating in MD will make the universe transit from the strong dissipative region to the weak dissipative region. Consequently, our numeric has confirmed our previous guess which the RD and MD are the weak dissipative system.